# MAROON: A Framework for the Joint Characterization of Near-Field High-Resolution Radar and Optical Depth Imaging Techniques

### v2

@@ Abstract.
Time of Flight, Radar Imaging, Radio Frequency, mmWave Imaging, MIMO Radar, Depth Camera, Spatial Calibration, Multimodal Sensor Fusion, Dataset
@@ 1. Introduction
Real-world computer-assisted tasks, for instance in robotics and tracking applications, frequently require the immediate assessment of spatial information to accurately reason about the environment at a specific point in time, which has led to the development of several single-view range and depth sensors. For autonomous driving, it has been shown that utilizing multimodal depth sensing techniques from both, the optical (lidar) and radio-frequency (radar) domain, can lead to superior performance and robustness in computer-assisted tasks (Velasco-Hernandez et al., 2020). Due to its environment, the autonomous driving industry has traditionally concentrated on far-field range sensing, with an unambiguous range of several meters and beyond. As recent high-resolution radio-frequency technologies utilize the concept of radar imaging to produce 3D information in form of a depth map — similar to optical depth or RGB-D cameras — they also become more popular in close range, where the target of interest is up to a few decimeters away from the sensor; however, a comprehensive and detailed characterization of these radar imaging technologies, which frequently operate in the radar’s near field, has yet to be realized.
@@
Drawing on prior research about wavelength-specific strengths and weaknesses, this paper addresses the unique challenge of characterizing various optical depth-imaging techniques alongside a high-resolution multiple-input multiple-output (MIMO) imaging radar in the near field. The latter is interchangeably referred to as a radio-frequency (RF) Time-of-Flight (ToF) depth imager. To this end, we mutually calibrated sensors of four different depth sensing technologies, that is active and passive stereo, near-infrared (NIR) ToF, and RF ToF in the millimeter-wave range.
Using these sensors, we captured the comprehensive MAROON dataset of various household objects and construction materials, of which example data is shown in Figure 2. The dataset enables an exploration of sensor-specific characteristics with respect to various object materials, geometries and distances from the sensor. Furthermore, the spatial calibration enables a direct comparison of all sensor reconstructions in a common metrical space. We then compare these results to a reference ground truth reconstruction from a state-of-the-art multi-view stereo (MVS) system. To achieve this, we developed a joint sensor evaluation framework that measures reconstruction differences between sensors and ground truth using different metrics, and that provides supplementary visualizations tailored to identify sensor-specific trends across multiple objects. Our findings reveal fundamental differences between optical and RF sensors, as well as characteristics arising from specific depth measurement techniques within a single frequency domain. We detail sensor-specific characteristics based on several key insights. First, by assessing deviation from the ground truth across different objects and object-to-sensor distances, we identify common sensor-specific trends. Given these, RF ToF reconstructions are found to be less complete compared to optical sensors. Lastly, we demonstrate that ToF sensors exhibit systematic depth errors resulting from interactions within partially transmissive media, such as subsurface scattering.
To summarize, our contributions are the following:
  • A framework enabling the systematic comparative characterization of depth-sensing techniques in both, the optical and radio-frequency signal domain.
  • A novel multimodal dataset, MAROON, comprising common objects in the near field, which were captured with the unique setup of four jointly calibrated depth sensors.
  • We publicly release this dataset together with the raw radar measurements, thus enabling the community to explore different signal reconstruction and filtering techniques. [^1] State-of-the-art methods for reconstruction and filtering will be publicly released together with the dataset viewer upon acceptance.
@@ 2.1. Optical Depth Sensing
Considering the sensor technologies, metrological research has been conducted in terms of optical ToF (Xiong et al., 2017; Zanuttigh et al., 2016) and active stereo (Giancola et al., 2018; Wang and Shih, 2021). Furthermore, working principles of passive stereo sensors have been widely addressed in computer vision algorithms (Szeliski, 2022). Similar to our work, Chiu et al. (2019) and Halmetschlager-Funek (2019) jointly characterize ToF and active stereo.
@@
Moreover, we discuss related frameworks for jointly characterizing sensors. Halmetschlager-Funek et al. (2019) compare individually estimated depth values against manual measurements. Chiu et al. (2019) and Giancola et al. (2018) align the 3D data captured from sensors with real or synthetic ground truth data, respectively. Most similar to ours, Hansard et al. (2012) analyze ToF and structured light sensors using a spatial calibration and investigated object material, color, geometry, and texture using ground-truth data obtained from a structured light scanner.
@@ 2.3. Fusion of RF and Optical Sensors in Close Range
Knowledge about complementary strengths is important for both, sensor characterization and sensor fusion. Research on multimodal sensor fusion in close range is very limited and mostly focused on capturing humans.
@@ 3. Preliminaries
[…]
mystyleleftmargin=0.0cm,rightmargin=0.0cm, innerleftmargin=0.1cm, innerrightmargin=0.1cm, innertopmargin=0.1cm, innerbottommargin=0.1cm {mdframed}[style=mystyle]Depth Imager. A sensor that, directly or indirectly, captures a depth image $\bm{D}$ of resolution $W\times H$, where each pixel $(u,v)$ contains a depth value $d$ measured along the axis perpendicular to the image plane. The depth may be indirectly measured from range and pixel position. We show the difference between range and depth in Figure 3.
Transmitter and Receiver. Optical receivers are small cells of image sensors, with a direct mapping to pixels. Transmitters are commonly LEDs or projectors. RF sensors have transmitting (TX) and receiving (RX) antennas.
Sensor. Describes all physical parts required for depth sensing and their spatial arrangement. Optical sensors typically consist of one or two cameras. Active sensors contain an additional illumination unit. RF sensors usually have one or more antenna arrays in different arrangements.
Depth Image Resolution $\mathbf{W\times H}$. The number of depth samples computed from the incoming signal. In cameras, the depth samples are directly computed for each pixel, i.e., each receiver. MIMO imaging radars apply signal post-processing to compute depth from the signal diversity at different receiver positions. Hence, the image resolution is not directly affected by the number of receivers but by the signal processing parameters such as the voxel density (cf. Section 4.2.2). While, in theory, the depth image resolution can be indefinitely high, in practice it is limited by the spatial resolution.
@@
Spatial Resolution $\bm{\delta}$. The term resolution has several definitions. Here, we explicitly refer to spatial resolution as the minimum distance between two points in space that can be resolved from the received signal. Lower spatial resolution means higher minimum resolvable distance, so more incorrect measurements are made when points become inseparable, as seen in Figure 3. Spatial resolution is a theoretical measure, and external factors such as the sensor design can affect its effective resolution. Literature about traditional multi-static RF ToF sensors divides spatial resolution into range resolution ($\delta_{r}$), and cross-range resolution along the horizontal ($\delta_{h}$) and vertical ($\delta_{v}$) axes, respectively. While $\delta_{h}$ and $\delta_{v}$ are measured at a specific range, a more general formulation is the angular resolution, often referred to as azimuth ($\omega_{h}$) and elevation ($\omega_{v}$) resolution (Willis and Griffiths, 2007). Literature from the optical domain shares a similar definition, however, with a different terminology and refers to depth resolution (Zanuttigh et al., 2016) ($\delta_{z}$) and pixel resolution ($\delta_{h}$, $\delta_{v}$) instead. Similar to traditional RF ToF sensors, optical sensors assume that the target is situated in the far field, where depth is approximately the same as range and, hence, $\delta_{z}\approx\delta_{r}$. We note that this assumption does not generally hold in the near field. Due to the concept of an expanded antenna aperture with multiple transmitters and receivers, the resolution of near-field MIMO imaging radars is defined with respect to the three orthogonal axes $x,y,z$. In these, $z$ refers to the depth axis and $x,y$ are parallel to the antenna aperture. Contrary to their difference in definition, they share the same terminology as far-field RF ToF sensors such that $\delta_{z}$ is referred to as range resolution and $\delta_{x,y}$ is the cross-range (Ahmed, 2014) or lateral (Ahmed, 2021) resolution, respectively. We illustrate the respective definitions that are used for optical ($\delta_{h}$, $\delta_{v}$, $\delta_{r}$) and RF sensors ($\delta_{x}$, $\delta_{y}$, $\delta_{z}$) in Figure 3. For simplicity, we define spatial resolution only for the sensor center, where depth and range are approximately the same, such that $\delta_{z}\approx\delta_{r}$ and $\delta_{x,y}\approx\delta_{h,v}$.
@@
Figure 3. Visualization of the effects caused by limited spatial resolution for multiple point targets. Optical sensors (left) have similar definitions as far-field RF ToF sensors and divide spatial resolution into depth $\delta_{r}\approx\delta_{z}$ and pixel resolution, $\delta_{h,v}$. Contrary to that, near-field imaging radars (right) refer to range $\delta_{z}$ and cross-range $\delta_{x,y}$ resolution. We define $\delta_{z}\approx\delta_{r}$ and $\delta_{x,y}\approx\delta_{h,v}$ for the sensor center, yet emphasize the conceptual difference between range and depth.
@@ 4.1. Characterization of Wavelength
Optical sensors utilize camera lenses to refract the received signal, which enable a precise focus onto nanometer-sized pixels and, at the same time, exhibit a high angular resolution. In the context of the mmWave domain, a camera analogue can be conceptualized as a single-input multiple-output (SIMO) radar, which is to say, a sensor comprising a single transmitter and multiple receivers. As indicated by Equation 1, mmWave sensors have a significantly lower angular resolution than optical sensors. Thus, high-resolution SIMO radars require comparably large antenna arrays with large lenses, which has proven to be impractical. Instead, high-resolution RF imaging sensors often are synthetic aperture radars (SAR), which use digital beamforming to focus. They utilize the angle diversity from distinct transmitter and receiver positions, which form a virtual aperture of size $L$, to increase the angular resolution (Bliss and Forsythe, 2003) and require fewer antennas compared to SIMO systems. The majority of near-field SAR radars is implemented with MIMO arrays, that is, with multiple transmitters and receivers.
@@ 4.2.1. Spatially Resolved Depth Sensing
Active stereo sensors assist correspondence finding with an illumination unit that projects a pattern onto the target, usually in the NIR range, captured by the two cameras. The signal-multiplexed (Zanuttigh et al., 2016) pattern assists epipolar correspondence matching in addition to shading and texture cues, improving depth quality in textureless regions and low light. However, challenges include pattern distortions and signal oversaturation at the NIR receiver in bright conditions. Further details on spatially resolved depth sensors are provided in the supplementary material.
4.2.2. Time-Resolved Sensors (Time of Flight)
Time of Flight is an active depth sensing method, in which depth is derived from the round-trip propagation time that it takes for a signal to be transmitted and received. The majority of ToF sensors utilized in the near field employ continuous wave (CW) signal modulations, which measure time based on the relative phase shift $\Delta\varphi$ between the transmitted and received signal. The depth is derived from the range $r$, which is measured by (Zanuttigh et al., 2016):
(2) $r=\text{c}\frac{\Delta\varphi}{2\pi f}\;.$
The signal frequency is denoted as $f$, and c is the speed of light in vacuum, which closely matches that of light in air. For further details about the operating principle, we refer to the supplementary material. ToF technologies employ a simplified model for range sensing, which assumes that targets are diffuse scatterers, with each signal reflecting directly from the first target. As a result, these technologies are sensitive to multi-path interference, where signals propagate indirectly in multiple directions before reaching the receiver. In Section 8, we identify partially transmissive materials as a major cause of such interference. Furthermore, the unambiguous range of resolving $\Delta\varphi$ is limited to the periodicity of the sinusoidal CW signal. To extend this range, the carrier signal can be modulated over time. In our experiments, we use ToF sensors with amplitude-modulated continuous wave (AMCW) and frequency-stepped continuous wave (FSCW) signal modulations, which are illustrated in Figure 4. Up next, we will discuss the operating principles of these depth sensing methods.
NIR AMCW Time of Flight
AMCW ToF algorithms usually operate on the SIMO principle, as they do not require as expensive sensor apertures as imaging radars, and often have more receivers and transmitters than can be effectively managed computationally in MIMO depth estimation algorithms (Zanuttigh et al., 2016). They modulate the amplitude $A$ of a carrier signal over time $t$ using a repetitive modulation signal $s_{\text{m}}$ such that the transmitted signal $s_{\text{t}}$ can be described as:
@@ 4.2.2. Time-Resolved Sensors (Time of Flight)
A constant phase offset is described by $\phi_{\text{c}}$. To extract the phase shift from the received signal $s_{\text{r}}$, it is cross-correlated with a so-called signal hypothesis $s_{\text{h}}$ (Zanuttigh et al., 2016):
[…]
$T_{\text{m}}$ is the period of the modulation signal $s_{\text{m}}$. Commonly, $s_{\text{h}}$ is chosen as the currently transmitted signal $s_{\text{t}}$ such that $c_{\text{r}}$ describes the signal similarity from which the relative phase shift to $s_{\text{r}}$ is inferred. Extracting this shift requires solving a multivariable equation system with parameters such as the received amplitude and external illumination. To achieve this, $c_{\text{r}}$ and, consequently, $s_{\text{r}}$ are commonly sampled at four points within $T_{\text{m}}$ (four-bucket-method) (Giancola et al., 2018). During the acquisition of those samples, AMCW ToF is affected by environmental changes, such as varying external NIR illumination and motion. Similar to active stereo, oversaturation of the NIR receiver may cause invalid signal responses.
MIMO FSCW Time of Flight
Given the frequency band $\text{b}=f_{\text{max}}-f_{\text{min}}$, FSCW ToF sensors iteratively send $\text{N}_{f}$ signals of one frequency in steps of $\Delta f=\text{b}/\text{N}_{f}$ (Bräunig et al., 2023b). More specifically, the transmitted signal $s_{\text{t}}$ of one capture can be described as:
@@
We denote the time window of one frequency step as $\Delta t$. Time-division multiplexing (TDM) avoids signal interference and facilitates the separation of the received signal into its originating transmitter and frequency components. SAR signal processing computes the depth $d$ and the pixel position $(u,v)$ from the phase shift and the angular diversity originating from multiple transmitting and receiving positions. For MIMO imaging radars, the state-of-the-art algorithm of backprojection (BP) (Ahmed, 2021) computes confidence values about a target’s presence in 3D space. This is achieved on the basis of local feature distributions within a volume based on the integrated signal of each RX-TX antenna pair. Similar to the previous section, the BP algorithm computes a correlation between the received signal $s_{\text{r}}$ and a signal hypothesis $s_{\text{h}}$:
(7) $c_{\text{BP}}(\underbrace{x,y,z}_{\bm{p}})=\sum_{n=1}^{\text{N}_{f}}\sum_{i=1}^{\text{N}_{\text{RX}}}\sum_{j=1}^{\text{N}_{\text{TX}}}s_{\text{r}}(f_{n},\bm{r}_{i},\bm{t}_{j})s_{\text{h}}(f_{n},\bm{r}_{i},\bm{t}_{j},\bm{p})\;.$
@@
A hypothesis is made on the basis of the transmitted signal, which is assumed to reflect at a point $\bm{p}\in\mathbb{R}^{3}$ in the sensor coordinate system, commonly sampled from a voxel grid of size $N_{v}=N_{x}\times N_{y}\times N_{z}$. The demodulated received signal, $s_{\text{r}}$, varies in transmit frequency $f_{n}=f_{\text{m}}(n)$, transmitter position $\bm{t}_{j}\in\mathbb{R}^{3}$, and receiver position $\bm{r}_{i}\in\mathbb{R}^{3}$. We further denote the numbers of transmitters and receivers as $\text{N}_{\text{TX}}$, and $\text{N}_{\text{RX}}$, respectively. Generally, hypotheses are made by assuming the signal propagation is following the so-called Born approximation (Ahmed, 2014). The result of the above equation is a complex phasor, $c_{\text{BP}}$, with its absolute value representing the confidence of a target’s presence, which is visualized in Figure 2. It is calculated from $s_{\text{r}}$ and $s_{\text{h}}$, which are analytic signals in complex notation. More details are given by Ahmed (2021). To compute a 2D depth map from the 3D voxel grid, an orthogonal maximum (intensity) projection (Bräunig et al., 2023b) is performed for each pixel $(u,v)=(x,y)$ along the cross-depth axes of the voxel grid:
[…]
@@
Besides projection, the confidence values are used as thresholds for depth filtering, that is, to distinguish target depth from sidelobes and background noise. The absolute value of the confidence $c_{\text{BP}}$ depends on both, the received phase and amplitude. Besides the depth, reasons for varying amplitude and phase over different object materials and geometries are manifold, and further insights will be given in Section 9. As a result, it is challenging to generalize the depth filtering process for unknown objects. Similar to NIR AMCW ToF, a MIMO FSCW ToF sensor requires multiple signal samples to compute depth and, thus, is sensitive to environmental changes during capture.
ZED X Mini (2.2 mm) (Stereolabs, 2023) Realsense D435i (Intel, 2023) Azure Kinect (Microsoft, 2022) QAR5 (Submodule) (Rohde & Schwarz, 2023)
@@
Depth Sensing Technology Passive Stereoscopy Active Stereoscopy Time of Flight (NIR) Time of Flight (RF)

### v4

# MAROON: A Dataset for the Joint Characterization of Near-Field High-Resolution Radio-Frequency and Optical Depth Imaging Techniques

@@ Abstract.
Time-of-Flight, Radar Imaging, Radio Frequency, mmWave Imaging, MIMO Radar, Depth Camera, Spatial Calibration, Multimodal Sensor Fusion, Dataset
@@ 1. Introduction
Real-world computer-assisted tasks, for instance in robotics and tracking applications, frequently require the immediate assessment of spatial information to accurately reason about the environment at a specific point in time, which has led to the development of several single-view range and depth sensors. For autonomous driving, it has been shown that utilizing multimodal depth sensing techniques from both the optical (lidar) and radio-frequency (radar) domain can lead to superior performance and robustness in computer-assisted tasks (Velasco-Hernandez et al., 2020). Due to its environment, the autonomous driving industry has traditionally concentrated on far-field range sensing, with an unambiguous range of several meters and beyond. As recent high-resolution radio-frequency technologies utilize the concept of radar imaging to produce 3D information in form of a depth map — similar to optical depth or RGB-D cameras — they also become more popular in close range, where the target of interest is up to a few decimeters away from the sensor; however, a comprehensive and detailed characterization of these radar imaging technologies, which frequently operate in the radar’s near field, is yet to be realized.
@@
Drawing on prior research about wavelength-specific strengths and weaknesses, this paper addresses the unique challenge of characterizing various optical depth-imaging techniques alongside a high-resolution multiple-input multiple-output (MIMO) imaging radar in the near field. The latter is interchangeably referred to as a radio-frequency (RF) Time-of-Flight (ToF) depth imager. To this end, we mutually calibrated sensors of four different depth sensing technologies, that is active and passive stereo, near-infrared (NIR) amplitude-modulated continuous wave (AMCW) ToF, and RF frequency-stepped continuous wave (FSCW) ToF in the millimeter-wave range.
There is a notable lack of multimodal datasets suitable for close-range applications, and, to our knowledge, this work is the first to incorporate imaging radars in this research area. With this in mind, we captured the MAROON dataset of various household objects and construction materials, of which example data is shown in Figure 2. Utilizing a high-resolution MIMO imaging radar, with a spatial resolution currently far beyond prevalent RF imaging sensors, we captured this dataset with a multitude of key objectives:
  1. (1)
    To evaluate sensor-specific reconstructions, considering various object materials, geometries, and distances to the sensors.
  2. (2)
    To establish a public data base for multimodal reconstruction research in close-range applications, bridging the radio-frequency and optical domains.
  3. (3)
    To characterize the under-researched effects of millimeter waves in near-field imaging radars, e.g. object materials in the radio-frequency domain, akin to studies on bi-directional reflectance distribution functions (BRDFs) in optics.
  4. (4)
    To improve radio-frequency signal simulations by supplying data for lower-resolution radar architectures and comparing synthetic signals with real measurements and a ground truth.
Together with the dataset, we developed a joint sensor evaluation framework that measures reconstruction differences between sensors and a ground truth using different metrics, providing supplementary visualizations tailored to identify sensor-specific trends across multiple objects. By analyzing these trends, we identified ToF scattering effects in partially transmissive materials and examined RF ToF reconstructions, which are typically less complete than those from optical sensors.
Moreover, we utilize the multimodal data of MAROON in two example applications: first, we show that the dataset serves as foundation for addressing inverse rendering problems in the radio-frequency domain. Taking up on concurrent work (Hofmann et al., 2025), we determine object-specific material properties, which are crucial in high-fidelity radar simulation Second, the variety of challenging objects in our dataset serves as a benchmark for developing novel multimodal reconstruction algorithms, as demonstrated in (Wirth et al., 2025). We extend this benchmark by additional experiments, varying the sensor configuration.
To summarize, our contributions include:
  • A novel multimodal dataset, MAROON, comprising common objects in the near field, captured using a jointly calibrated setup of three optical depth sensors, a high-resolution imaging radar, and high-quality multi-view reconstruction for ground-truth geometry. We are releasing this dataset alongside the raw radar measurements to facilitate exploration of various signal reconstruction and filtering techniques.
@@
  • Two applications of the dataset: inverse rendering for material characterization and high-speed multimodal reconstruction.
@@ 2.1. Optical Depth Sensing
Considering the sensor technologies, metrological research has been conducted in terms of optical ToF (Xiong et al., 2017; Zanuttigh et al., 2016) and active stereo (Giancola et al., 2018; Wang and Shih, 2021). Furthermore, working principles of passive stereo sensors have been widely addressed in computer vision algorithms (Szeliski, 2022). Similar to our work, Chiu et al. (2019) and Halmetschlager-Funek et al. (2019) jointly characterize ToF and active stereo.
@@
Moreover, we discuss related frameworks for jointly characterizing sensors. Halmetschlager-Funek et al. (2019) compare individually estimated depth values against manual measurements. Chiu et al. (2019) and Giancola et al. (2018) align the 3D data captured from sensors with real or synthetic ground-truth data, respectively. Most similar to ours, Hansard et al. (2012) analyze ToF and structured light sensors using a spatial calibration and investigated object material, color, geometry, and texture using ground-truth data obtained from a structured light scanner.
@@ 2.3. Fusion of RF and Optical Sensors in Close Range
Knowledge about complementary strengths is important for both, sensor characterization and sensor fusion. While significant research efforts have been devoted to the field of autonomous driving — where radar sensors primarily operate in their respective far field — research on multimodal sensor fusion in close range is very limited and mostly focused on capturing humans.
@@ 3. Preliminaries
Depth Imager. A sensor that, directly or indirectly, captures a depth image $\bm{D}$ of resolution $W\times H$, where each pixel $(u,v)$ contains a depth value $d$ measured along the axis perpendicular to the image plane. The depth may be indirectly measured from range and pixel position. We show the difference between range and depth in Figure 3. Transmitter and Receiver. Optical receivers are small cells of image sensors, with a direct mapping to pixels. Transmitters are commonly LEDs or projectors. RF sensors have transmitting (TX) and receiving (RX) antennas. Sensor. Describes all physical parts required for depth sensing and their spatial arrangement. Optical sensors typically consist of one or two cameras. Active sensors contain an additional illumination unit. RF sensors usually have one or more antenna arrays in different arrangements. Depth Image Resolution $\mathbf{W\times H}$. The number of depth samples computed from the incoming signal. In cameras, the depth samples are directly computed for each pixel, i.e., each receiver. MIMO imaging radars apply signal post-processing to compute depth from the signal diversity at different receiver positions. Hence, the image resolution is not directly affected by the number of receivers but by the signal processing parameters such as the voxel density (cf. Section 4.2.2). While, in theory, the depth image resolution can be indefinitely high, in practice it is limited by the spatial resolution. Spatial Resolution $\bm{\delta}$. The term resolution has several definitions. Here, we explicitly refer to spatial resolution as the minimum distance between two points in space that can be resolved from the received signal. Lower spatial resolution means higher minimum resolvable distance, so more incorrect measurements are made when points become inseparable, as seen in Figure 3. Spatial resolution is a theoretical measure, and external factors such as the sensor design can affect its effective resolution. Literature about traditional multi-static RF ToF sensors divides spatial resolution into range resolution ($\delta_{r}$), and cross-range resolution along the horizontal ($\delta_{h}$) and vertical ($\delta_{v}$) axes, respectively. While $\delta_{h}$ and $\delta_{v}$ are measured at a specific range, a more general formulation is the angular resolution, often referred to as azimuth ($\omega_{h}$) and elevation ($\omega_{v}$) resolution (Willis and Griffiths, 2007). Literature from the optical domain shares a similar definition, however, with a different terminology and refers to depth resolution (Zanuttigh et al., 2016) ($\delta_{z}$) and pixel resolution ($\delta_{h}$, $\delta_{v}$) instead. Similar to traditional RF ToF sensors, optical sensors postulate that the target is situated in the far field, where depth is assumed to be approximately the same as range and, hence, $\delta_{z}\approx\delta_{r}$. We note that this assumption is not accurate in the near field. Due to the concept of an expanded antenna aperture with multiple transmitters and receivers, the resolution of near-field MIMO imaging radars is defined with respect to the three orthogonal axes $x,y,z$. In these, $z$ refers to the depth axis and $x,y$ are parallel to the antenna aperture. Contrary to their difference in definition, they share the same terminology as far-field RF ToF sensors such that $\delta_{z}$ is referred to as range resolution and $\delta_{x,y}$ is the cross-range (Ahmed, 2014) or lateral (Ahmed, 2021) resolution, respectively. We illustrate the respective definitions that are used for optical ($\delta_{h}$, $\delta_{v}$, $\delta_{r}$) and RF sensors ($\delta_{x}$, $\delta_{y}$, $\delta_{z}$) in Figure 3. For simplicity, we define spatial resolution only for the sensor center, where depth and range are approximately the same, such that $\delta_{z}\approx\delta_{r}$ and $\delta_{x,y}\approx\delta_{h,v}$.
@@
Figure 3. Visualization of the effects caused by limited spatial resolution for multiple point targets. Optical sensors (left) have similar definitions as far-field RF ToF sensors and divide spatial resolution into depth, $\delta_{r}$, and pixel resolution, $\delta_{h,v}$. Contrary to that, near-field imaging radars (right) refer to range $\delta_{z}$ and cross-range $\delta_{x,y}$ resolution. We assume $\delta_{z}\approx\delta_{r}$ and $\delta_{x,y}\approx\delta_{h,v}$ for the sensor center, yet emphasize the conceptual difference between range and depth.
@@ 4.1. Characterization of Wavelength
Optical sensors utilize camera lenses to refract the received signal, which enable a precise focus onto nanometer-sized pixels and, at the same time, exhibit a high angular resolution. In the context of the mmWave domain, a camera analogue can be conceptualized as a single-input multiple-output (SIMO) radar, i.e., a sensor comprising a single transmitter and multiple receivers. As indicated by Equation 1, mmWave sensors have a considerably lower angular resolution than optical sensors. Thus, high-resolution SIMO radars require comparably large antenna arrays with large lenses, which has proven to be impractical. Instead, high-resolution RF imaging sensors often are synthetic aperture radars (SAR), which use digital beamforming to focus. They utilize the angle diversity from distinct transmitter and receiver positions, which form a virtual aperture of size $L$, to increase the angular resolution (Bliss and Forsythe, 2003) and require fewer antennas compared to SIMO systems. The majority of near-field SAR radars is implemented with MIMO arrays, that is, with multiple transmitters and receivers.
@@ 4.2.1. Spatially Resolved Depth Sensing
Active stereo sensors assist correspondence finding with an illumination unit that projects a pattern onto the target, usually in the NIR range, captured by the two cameras. The signal-multiplexed (Zanuttigh et al., 2016) pattern supports epipolar correspondence matching in addition to shading and texture cues, improving depth quality in textureless regions and low light. However, challenges include pattern distortions and signal oversaturation at the NIR receiver in bright conditions. Further details on spatially resolved depth sensors are provided in the supplementary material.
4.2.2. Time-Resolved Sensors (Time-of-Flight)
Time-of-Flight is an active depth sensing method, in which depth is derived from the round-trip propagation time that it takes for a signal to be transmitted and received. The majority of ToF sensors utilized in the near field employ continuous wave (CW) signal modulations, which measure time based on the relative phase shift $\Delta\varphi$ between the transmitted and received signal. The depth is derived from the range $r$, which is measured by (Zanuttigh et al., 2016):
(2) $r=\text{c}\frac{\Delta\varphi}{4\pi f}\;.$
The signal frequency is denoted as $f$, and c is the speed of light in vacuum, which closely matches that of light in air. For further details about the operating principle, we refer to the supplementary material. ToF technologies employ a simplified model for range sensing, which assumes that targets are weak scatterers (Ahmed, 2014), with each signal reflecting directly from the first target. As a result, these technologies are sensitive to multi-path interference. In Section 7.2, we identify partially transmissive materials as a major cause of such interference. Furthermore, the unambiguous range, in which $\Delta\varphi$ can be correctly resolved, is limited to the periodicity of the sinusoidal CW signal. To extend this range, the carrier signal can be modulated over time. Noting that various modulation schemes exist, e.g., frequency-modulated continuous wave (FMCW) modulation, we use ToF sensors with AMCW and FSCW signal modulations, which are illustrated in Figure 4. Up next, we will discuss the operating principles of these depth sensing methods.
NIR AMCW Time-of-Flight
AMCW ToF algorithms modulate the amplitude $A$ of a carrier signal over time $t$ using a repetitive modulation signal $s_{\text{m}}$ such that the transmitted signal $s_{\text{t}}$ can be described as:
@@ 4.2.2. Time-Resolved Sensors (Time-of-Flight)
A constant phase offset is described by $\phi_{\text{c}}$. To extract the phase shift from the received signal, it is demodulated to yield $m_{\text{r}}$ and cross-correlated with a so-called signal hypothesis $s_{\text{h}}$ (Zanuttigh et al., 2016):
(4) $c_{\text{r}}(t)=\int_{0}^{T_{\text{m}}}{}{m_{\text{r}}}{}(t)s_{\text{h}}(t+t^{\prime})dt^{\prime}\;.$
$T_{\text{m}}$ is the period of the modulation signal. Commonly, $s_{\text{h}}$ is chosen as the currently transmitted signal $s_{\text{t}}$ such that $c_{\text{r}}$ describes the signal similarity from which the relative phase shift to $m_{\text{r}}$ is inferred. Extracting this shift requires solving a multivariable equation system with parameters such as the received amplitude and external illumination. To achieve this, $c_{\text{r}}$ and, consequently, $m_{\text{r}}$ are commonly sampled at four points within $T_{\text{m}}$ (four-bucket-method) (Giancola et al., 2018). During the acquisition of those samples, AMCW ToF is affected by environmental changes, such as varying external NIR illumination and motion. Moreover, oversaturation of the NIR receiver may cause invalid signal responses.
MIMO FSCW Time-of-Flight
FSCW ToF sensors model the frequency of the transmitting signal as a function of time. Given the frequency band $\text{b}=f_{\text{max}}-f_{\text{min}}$, they iteratively send $\text{N}_{f}$ signals of one frequency in steps of $\Delta f=\text{b}/(\text{N}_{f}-1)$ (Bräunig et al., 2023). More specifically, the transmitted signal $s_{\text{t}}$ of one capture can be described as:
@@
We denote the time window of one frequency step as $\Delta t$. Time-division multiplexing (TDM) avoids signal interference and facilitates the separation of the received signal into its originating transmitter and frequency components. SAR signal processing computes the depth $d$ and the pixel position $(u,v)$ from the phase shift and the angular diversity originating from multiple transmitting and receiving positions. For MIMO imaging radars, the state-of-the-art algorithm of backprojection (BP) (Wolf, 1969; Ahmed, 2021) computes confidence values about a target’s presence in 3D space. This is achieved on the basis of local feature distributions within a volume based on the integrated signal of each RX-TX antenna pair. Similar to the four-bucket-method for AMCW ToF, the BP algorithm computes a correlation between the received demodulated signal $m_{\text{r}}$ and a signal hypothesis $s_{\text{h}}$:
(7) $c_{\text{BP}}(\underbrace{x,y,z}_{\bm{p}})=\sum_{n=1}^{\text{N}_{f}}\sum_{i=1}^{\text{N}_{\text{RX}}}\sum_{j=1}^{\text{N}_{\text{TX}}}{}{m_{\text{r}}}{}(f_{n},\bm{r}_{i},\bm{t}_{j})s_{\text{h}}(f_{n},\bm{r}_{i},\bm{t}_{j},\bm{p})\;.$
@@
A hypothesis is made on the basis of the transmitted signal, which is assumed to reflect at a point $\bm{p}\in\mathbb{R}^{3}$ in the sensor coordinate system, commonly sampled from a voxel grid of size $N_{v}=N_{x}\times N_{y}\times N_{z}$. The demodulated received signal, $m_{\text{r}}$, varies in transmit frequency $f_{n}=f_{\text{m}}(n)$, transmitter position $\bm{t}_{j}\in\mathbb{R}^{3}$, and receiver position $\bm{r}_{i}\in\mathbb{R}^{3}$. The numbers of transmitters and receivers are denoted as $\text{N}_{\text{TX}}$, and $\text{N}_{\text{RX}}$, respectively. Generally, hypotheses are made by assuming the signal propagation is following the Born approximation (Ahmed, 2014). The result of the above equation is a complex phasor $c_{\text{BP}}$, calculated from $m_{\text{r}}$ and $s_{\text{h}}$, which are analytic signals in complex notation. To compute a 2D depth map from the 3D voxel grid, an orthogonal maximum (intensity) projection (Bräunig et al., 2023) is performed for each pixel $(u,v)=(x,y)$ along the cross-depth axes of the voxel grid:
(8) $d(u,v)=\underset{z}{\text{argmax}}\lVert c_{\text{BP}}(x,y,z)\rVert_{2}=\underset{z}{\text{argmax}}\ \kappa(x,y,z)\;.$
@@
The letter $\kappa$ denotes the so-called confidence of a target’s presence, as visualized in Figure 2. Besides projection, the confidence values are used as thresholds for depth filtering, that is, to distinguish target depth from sidelobes and background noise. As $\kappa$ directly relates to $c_{\text{BP}}$, it depends on both, the received phase and amplitude. Besides the depth, reasons for varying amplitude and phase over different object materials and geometries are manifold, and further insights will be given in Section 7.3. As a result, it is challenging to generalize the depth filtering process for unknown objects. Conversely, if a point $\bm{p}$ on the target is partially occluded for multiple RX-TX antenna pairs, resulting in a decrease in its confidence value — potentially to the level of background noise — it may be filtered out.
Similar to NIR AMCW ToF, a MIMO FSCW ToF sensor is sensitive to environmental changes while capturing multiple signal samples.
ZED X Mini (2.2 mm)
(Stereolabs, 2023)
Realsense D435i
(Intel, 2023)
Azure Kinect
(Microsoft, 2022)
QAR50 (Submodule)
(Rohde & Schwarz, 2023)
@@
Depth Sensing Technology Passive Stereoscopy Active Stereoscopy Time-of-Flight (NIR) Time-of-Flight (RF)